# LEMMA-RCA: A Large Multi-modal Multi-domain Dataset for Root Cause Analysis

### v1

@@ 4.2 Offline Root Cause Analysis Results
 Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10  Metric Only Dynotears 0 0 50.0% 6.96% 0 0 7.50% PC 0 0 25.0% 5.32% 0 0 5.0% C-LSTM 25.0% 75.0% 75.0% 47.39% 50.0% 25.0% 67.50% GOLEM 0 0 25.0% 4.31% 0 0 2.50% REASON 75.0% 100.0% 100.0% 87.50% 91.67% 95.0% 97.5% Log Only Dynotears 0 0 25.0% 5.84% 0 0 7.50% PC 0 0 25.0% 6.93% 0 0 12.50% C-LSTM 0 0 25.0% 5.90% 0 0 7.50% GOLEM 0 0 25.0% 5.83% 0 0 7.50% REASON 0 50.0% 75.0% 21.56% 16.67% 25.0% 40.0% Multi-Modality Dynotears 0 0 50.0% 9.52% 0 0 15.0% PC 0 0 25.0% 6.43% 0 0 12.5% C-LSTM 50.0% 75.0% 75.0% 59.25% 58.33% 65.0% 70.0% GOLEM 0 0 25.0% 6.43% 0 0 5.0% REASON 75.0% 100.0% 100.0% 87.50% 91.67% 95.0% 97.5% Nezha 0 50.0% 75.0% 19.29% 8.33% 25.0% 47.50% MULAN 100.0% 100.0% 100.0% 100.0% 100.0% 100.0% 100.0%
Table 5: Results for offline RCA baselines on the SWaT and WADI dataset.
 Dataset Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10  SWaT Dynotears 12.5% 32.29% 42.71% 27.85% 20.14% 24.38% 30.83% PC 12.5% 34.38% 58.33% 26.16% 12.85% 20.42% 35.0% C-LSTM 12.5% 28.13% 52.08% 29.35% 13.89% 17.71% 31.88% GOLEM 6.25% 12.5% 47.92% 22.36% 7.64% 9.58% 25.0% REASON 25.0% 66.67% 84.38% 40.99% 23.96% 35.0% 57.60%  WADI Dynotears 7.14% 30.0% 47.62% 22.23% 10.71% 17.43% 26.81% PC 7.14% 35.0% 50.0% 27.74% 16.27% 23.90% 34.57% C-LSTM 0% 35.0% 51.19% 24.40% 11.51% 18.55% 32.73% GOLEM 0% 40.0% 53.57% 23.48% 9.92% 20.38% 34.83% REASON 28.57% 65.0% 79.76% 53.35% 42.46% 50.62% 63.76%
Water Treatment/Distribution. We employ five single-modal RCA methods to assess root cause localization performance on the SWaT and WADI sub-datasets. The comparative results, presented in Table Table 5, are evaluated in terms of PR@K, MRR, and MAP@K. Consistent with observations on the Product Review and Cloud Computing sub-datasets, REASON outperforms the other four baseline methods. However, a decline in performance for the best baseline method, REASON, is noted when compared to its results on the Product Review and Cloud Computing datasets. This decrease in performance can be attributed to the nature of the SWaT and WADI sub-datasets, where faults are brief and the intervals between them are short. These fleeting events can be easily missed by most RCA methods, thus posing a significant challenge in accurately identifying the root causes within these two sub-datasets.
4.3 Online Root Cause Analysis Results
We evaluate three RCA methods on all sub-datasets to assess the quality of the LEMMA-RCA sub-dataset in an online setting. Notice that due to the lack of multi-modal online RCA methods, we measure the performance of these single-modal baseline methods using only metric data shown in Table 6. By observation, we find that the online version of RCA models (e.g., GOLEM) outperform their offline version (e.g., GOLEM) as online methods can rapidly capture the changing patterns of the metric data, thus learning a more accurate and noise-free causal structure for RCA. Among online methods, CORAL significantly outperforms NOTEARS and GOLEM due to the design of state-invariant and state-dependent representations learning tailored for the online setting. Notably, LEMMA-RCA is a large-scale real-world dataset, consisting of more than 100,000 timestamps across several days with various system fault scenarios, which can be naturally transformed to the online setting, compared with small datasets (e.g., NeZha [12]) with limited timestamps for online RCA.
Table 6: Results for online root cause analysis baselines on all sub-datasets.
 Dataset Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10   Product Review CORAL 75.0% 100.0% 100.0% 87.50% 91.67% 95.0% 97.50% NOTEARS 25.0% 75.0% 75.0% 48.11% 50.0% 60.0% 67.50% GOLEM 50.0% 75.0% 75.0% 64.58% 66.67% 70.0% 72.50%  SWaT CORAL 6.25% 55.21% 92.71% 31.72% 15.63% 29.79% 53.96% NOTEARS 6.25% 36.46% 67.71% 26.30% 14.93% 23.54% 42.19% GOLEM 6.25% 42.71% 68.75% 28.09% 17.01% 26.04% 43.65%  WADI CORAL 35.71% 60.0% 83.33% 51.90% 28.71% 36.05% 56.0% NOTEARS 14.29% 45.71% 72.62% 37.74% 18.65% 27.48% 48.38% GOLEM 21.43% 60.0% 73.81% 40.24% 19.84% 30.33% 48.98%
@@ 5 Discussions
Broader impact: To facilitate accurate, efficient, and multi-modal root cause analysis research across diverse domains, we introduce LEMMA-RCA as a new benchmark dataset. Our dataset also offers significant potential for advancing research in areas like multi-modal anomaly detection, change point detection, causal structure learning, and LLM-based system diagnosis. Based on the thorough data analysis and extensive experimental results, we highlight the following areas for future research:
@@
  • Online Multi-Modal Root Cause Analysis: The majority of RCA methods are offline and typically single-modal. However, our research highlights a notable gap in online, multi-modal approaches. This presents an opportunity to develop real-time, multi-modal root cause analysis methods that can instantly process and analyze diverse data streams. Such advancements are crucial for dynamic environments like industrial automation and real-time monitoring services, where immediate response is essential.
Limitations: Despite its broad capabilities, the LEMMA-RCA dataset may have limitations in terms of the dataset generalizability, as the system fault scenarios we created may not adequately reflect the diversity of conditions prevalent in broader real-world applications. Another limitation is the LEMMA-RCA dataset is associated with some missing data in the system metric readings. These issues can arise from factors such as system interruptions and other unforeseen circumstances.
@@ 6 Conclusion
In this work, we present LEMMA-RCA, the first large-scale, open-source dataset featuring real system faults across various application domains and multiple modalities. We evaluate the quality of LEMMA-RCA by testing the performance of eight baseline methodologies on this dataset under different settings, including offline/online modes and single/multiple-modality data. Our experimental results demonstrate the high quality of LEMMA-RCA. By making this dataset publicly available, we aim to facilitate further research and innovation in root cause analysis for complex systems, contributing significantly to the development of more robust and secure methodologies that ensure the high performance of modern systems, particularly those that are mission-critical.

### v3

@@ 4.1 Experimental Setup
 Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10  Metric Only PC 0 0 0.250 0.053 0 0 0.050 RCD 0 0.250 0.750 0.185 0.167 0.200 0.350 $\epsilon$-Diagnosis 0 0 0.250 0.038 0 0 0.050 CIRCA 0 0.750 0.750 0.283 0.250 0.450 0.600 BARO 0.250 0.250 0.250 0.286 0.250 0.250 0.250 Log Only PC 0 0 0.250 0.069 0 0 0.125 RCD 0 0 0.250 0.056 0 0 0.025 $\epsilon$-Diagnosis 0 0 0.250 0.038 0 0 0.050 CIRCA 0 0 0.500 0.080 0 0 0.125 BARO 0 0.250 0.250 0.139 0.167 0.200 0.225 Multi-Modality PC 0 0 0.250 0.064 0 0 0.125 RCD 0 0.500 0.750 0.231 0.167 0.300 0.525 $\epsilon$-Diagnosis 0 0 0.250 0.041 0 0 0.075 CIRCA 0 0.750 1.000 0.299 0.250 0.450 0.650 BARO 0.750 0.750 1.000 0.775 0.750 0.750 0.775 Nezha 0 0.500 0.750 0.193 0.083 0.250 0.475
4.2 Root Cause Analysis Results
Product Review and Cloud Computing. We evaluate six RCA methods including both single-modal and multi-modal methods on Product Review and Cloud Computing sub-datasets. The experimental results are presented in Table 3 and Table 4 with respect to Precision at K (PR@K), Mean Reciprocal Rank (MRR), and Mean Average Precision at K (MAP@K). Our observations reveal the following insights: (1) The PC algorithm and $\epsilon$-Diagnosis perform worst on both the Product Review and Cloud Computing sub-datasets. We conjecture that PC, RCD, and $\epsilon$-Diagnosis struggle to capture long-term dependencies in large-scale datasets, making it difficult to detect abnormal temporal patterns. (2) CIRCA outperforms RCD and $\epsilon$-Diagnosis, consistent with findings from the Petshop study [12], where CIRCA’s regression-based hypothesis testing and adjustment mechanisms led to higher diagnostic accuracy. (3) Multi-modal input—combining both metric and log data—significantly enhances the performance of RCA methods compared to using either modality alone. For example, BARO achieves only 25% PR@1 with metric data and 0% with log data on the Product Review sub-dataset, but reaches 75% PR@1 when both are used, correctly identifying the root cause in 75% of fault scenarios. This highlights the complementary nature of log and metric data and the importance of integrating both to improve diagnostic accuracy and overall performance, particularly in terms of MRR.
Table 4: Results for RCA with multiple modalities on the Cloud Computing sub-dataset.
 Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Metric Only PC 0 0 0 0.029 0 0 0 RCD 0 0.250 0.250 0.126 0.083 0.150 0.200 $\epsilon$-Diagnosis 0 0 0.250 0.067 0 0 0.025 CIRCA 0 0 0.500 0.105 0 0 0.150 BARO 0 0.250 0.250 0.105 0 0.100 0.175 Log Only PC 0 0 0 0.032 0 0 0 RCD 0 0 0 0.059 0 0 0 $\epsilon$-Diagnosis 0 0 0 0.059 0 0 0 CIRCA 0 0.250 0.250 0.110 0 0.100 0.175 BARO 0 0.250 0.250 0.105 0 0.100 0.175 Multi-Modality PC 0 0 0.250 0.064 0 0 0.125 RCD 0 0.250 0.250 0.092 0 0.050 0.150 $\epsilon$-Diagnosis 0 0 0.250 0.067 0 0 0.025 CIRCA 0 0.250 0.500 0.147 0.083 0.150 0.225 BARO 0 0.250 0.250 0.126 0.083 0.150 0.200 Nezha 0 0.250 0.250 0.105 0 0.100 0.175
Table 5: Results for RCA baselines on the SWaT sub-dataset.
 Dataset Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10  SWaT PC 0.125 0.344 0.583 0.262 0.129 0.204 0.350 RCD 0.125 0.125 0.625 0.228 0.125 0.125 0.344 $\varepsilon$-Diagnosis 0.125 0.125 0.563 0.217 0.125 0.125 0.294 CIRCA 0.188 0.250 0.688 0.287 0.188 0.200 0.394 BARO 0 0.208 0.208 0.124 0.083 0.133 0.171
Table 6: Results for RCA baselines on the WADI sub-dataset.
 Dataset Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10  WADI PC 0.071 0.350 0.500 0.277 0.163 0.239 0.346 RCD 0 0 .071 0.119 0.054 0.048 0.057 0.076 $\varepsilon$-Diagnosis 0 0 0.022 0.020 0 0 0.009 CIRCA 0.143 0.550 0.714 0.350 0.301 0.400 0.529 BARO 0 0.143 0.143 0.085 0.071 0.100 0.121
Water Treatment/Distribution. We evaluate five single-modal RCA methods on the SWaT and WADI sub-datasets using the same set of evaluation metrics. Table 5 presents the results for SWaT, while Table 6 shows the results for WADI. We observe that CIRCA consistently achieves the best overall performance on both datasets, although the PC algorithm occasionally outperforms it on SWaT by a small margin. This trend aligns with the observations from the Product Review and Cloud Computing sub-datasets. Notably, the results also indicate considerable room for improvement, likely due to the characteristics of the SWaT and WADI datasets—faults are short-lived, and the intervals between them are brief. These fleeting events are easily overlooked by most RCA methods, posing a significant challenge for accurate root cause identification.
@@ 5 Discussions
Broader impact: To facilitate accurate, efficient, and multi-modal root cause analysis research across diverse domains, we introduce LEMMA-RCA as a new benchmark dataset. Our dataset also offers significant potential for advancing research in areas like multi-modal anomaly detection, change point detection, and system diagnosis. Based on the thorough data analysis and extensive experimental results, we highlight the following areas for future research:
@@
  • Online and/or Multi-Modal Root Cause Analysis: Most RCA methods are offline and single-modal, leaving a gap for real-time and/or multi-modal approaches. Developing these methods can enable instant analysis of diverse data streams, essential for dynamic environments like industrial automation and real-time monitoring.
  • LLM-Based Root Cause Analysis: The emergence of LLMs presents new opportunities for RCA by enabling systems to reason over complex, unstructured, and heterogeneous data sources. Future research can explore how LLMs can be adapted or fine-tuned for root cause inference, how they can incorporate domain knowledge, and how their interpretability and reliability can be improved in safety-critical or high-stakes environments.
Limitations: Despite its broad capabilities, the LEMMA-RCA dataset may have limited generalizability, as its fault scenarios may not fully capture the diversity of real-world conditions due to factors like system interruptions and unforeseen circumstances.
@@ 6 Conclusion
In this work, we introduce LEMMA-RCA, the first large-scale, open-source dataset featuring real system faults across multiple application domains and data modalities. We conduct a comprehensive empirical study using six baseline RCA methods, evaluating their performance on both single-modal and multi-modal data. The experimental results highlight the utility of LEMMA-RCA as a benchmarking resource. By releasing this dataset publicly, we aim to advance research in root cause analysis for complex systems and support the development of more robust and reliable methodologies, particularly for mission-critical applications.
Appendix A Monitoring Time Series Segmentation for SWaT and WADI
In the original SWaT and WADI datasets, the attack model demonstrates irregular attack patterns, occasionally targeting multiple sensors simultaneously, or executing attacks at closely spaced intervals. To follow the principles of RCA, we have established two specific preprocessing rules for these datasets: 1) Each recorded attack event must only involve a single sensor or actuator. 2) The duration of the dataset corresponding to each attack event must be standardized to two hours. Consequently, we selectively keep attack events that impact only one sensor or actuator. If the interval between successive attack events is insufficiently short, we assume the stability in the monitoring data immediately before and after each attack event. To ensure the necessary two-hour duration for each event, we concatenate normal-state data from both before and after the attack period. This adjustment positions the attack event centrally within a continuous two-hour segment, facilitating consistent and accurate analysis.
Appendix B Additional System Fault Scenarios
This section describes the processes used to generate and monitor system fault scenarios, with emphasis on mimicking real-world fault patterns. Each scenario involved the induction of specific failure conditions, while allowing the microservice system to exhibit its natural behavior under stress. Metrics and logs were collected using established monitoring tools, such as Prometheus, Elasticsearch, CloudWatch, Jaeger, and JMeter.
  • Silent Pod Degradation Fault.
    • Description: A pod in a load balancer contains a latent bug causing its CPU usage to rise, which gradually increases latency for a subset of users without triggering autoscaling or error alerts.
    • Method: We periodically sent requests to Microservice A over a 24-hour period. After this initial observation, we manually increased the CPU load on one specific productpage-v1 pod to simulate the bug.
    • Data Collection: Metrics and logs were collected from CloudWatch, while KPIs such as latency were measured using JMeter. The goal was to trace latency increases back to the specific pod with elevated CPU utilization.
  • Noisy Neighbor Issue.
    • Description: A neighboring pod in a shared node generates high CPU load, impacting the performance of the productpage-v1 pod and causing elevated error rates.
    • Method: Requests were sent to Microservice A, while the pod ratings of Microservice B (robot-shop) were moved to the same node as productpage-v1, generating contention.
    • Data Collection: Metrics (CPU usage, memory usage) were gathered using Prometheus, while logs were obtained from CloudWatch Logs. Configuration changes, such as node assignments, were also recorded.
  • Node Resource Contention Stress Test.
    • Description: A stress test on CPU resources was conducted by inducing high load on Microservice B, co-located with Microservice A on the same node.
    • Method: Periodic requests were sent to Microservice A using JMeter, while a high CPU load was generated on Microservice B using the OpenSSL speed command.
    • Data Collection: HTTP response logs from JMeter were analyzed for performance impacts. System metrics (CPU and memory usage) were retrieved from Prometheus, while container logs were collected from Elasticsearch.
  • DDoS Attack.
    • Description: A Distributed Denial of Service (DDoS) attack was simulated to overload the system, causing Out-of-Memory (OOM) errors in targeted pods.
    • Method: Over a monitoring period of approximately 48 hours, we gradually increased the request rate to Microservice A, eventually overwhelming the reviews-v2 and reviews-v3 pods.
    • Data Collection: Metrics such as CPU and memory utilization were collected via Prometheus. Logs from Jaeger and Elasticsearch provided insights into the system’s response to the attack.
  • Malware Attack.
    • Description: A malware pod executed a password list attack to compromise other pods, propagating DDoS scripts to degrade overall system performance.
    • Method: The attack started from a designated pod (scenario10-malware-deployment) and targeted others via SSH password brute-forcing, ultimately generating high load on productpage-v1.
    • Data Collection: JMeter was used to monitor KPIs (latency, error rate), while Prometheus and CloudWatch Logs provided system metrics and logs for root-cause analysis.
  • Bug Infection.
    • Description: A latent bug in the API caused asymmetric CPU load increases, degrading response times without fully utilizing the CPU capacity.
    • Method: Requests were sent periodically to the web service, and after a day, a script induced increased CPU utilization on one core.
    • Data Collection: KPIs were measured using JMeter, while system metrics and logs were collected via CloudWatch for detailed analysis.
  • Configuration Fault.
    • Description: An incorrect resource limit in a Kubernetes manifest file led to a pod being terminated by the OOM killer, impacting other services.
    • Method: Requests were sent to Microservice A, while a Git push introduced a faulty configuration for the details-v1 pod. The misconfigured pod eventually failed under load.
    • Data Collection: Error rates were tracked using JMeter, and metrics/logs were retrieved from Prometheus and CloudWatch for root-cause identification.
Appendix C LEMMA-RCA License
The LEMMA-RCA benchmark dataset is released under a CC BY-ND 4.0 International License: https://creativecommons.org/licenses/by-nd/4.0. The license of any specific baseline methods used in our codebase should be verified on their official repositories.
Appendix D Reproducibility
All experiments are conducted on a server running Ubuntu 18 with an Intel(R) Xeon(R) Silver 4110 CPU @2.10GHz and one 11GB GTX2080 GPU. In addition, all methods were implemented using Python 3.8.12 and PyTorch 1.7.1.
Appendix E Detailed Description of Baselines
We evaluate the performance of the following RCA models on the benchmark sub-datasets:
  • PC-based [35]: It is an unsupervised root cause analysis method designed for complex systems using multiple types of performance metrics. It constructs an anomaly behavior graph using a PC-based causal discovery algorithm to capture service-level dependencies. Based on this graph, it applies heuristic random walk strategies—including forward, backward, and self-directed walks—to trace and rank potential root causes.
  • CIRCA [37]: CIRCA is an unsupervised root cause analysis method that formulates the problem as a causal inference task called intervention recognition. Its core idea is to identify root cause indicators by evaluating changes in the probability distribution of monitoring variables conditioned on their parents in a Causal Bayesian Network (CBN). CIRCA applies this approach to online service systems by constructing a graph among monitoring metrics, leveraging system architecture knowledge and causal assumptions to guide the analysis.
  • $\epsilon$-Diagnosis [38]: $\epsilon$-Diagnosis is an unsupervised, low-cost diagnosis algorithm designed to address small-window long-tail latency (SWLT) in web services, which arises in short statistical windows and typically affects a small subset of containers in microservice clusters. It uses a two-sample test algorithm and $\epsilon$-statistics to measure the similarity of time series, enabling the identification of root-cause metrics from millions of metrics. The algorithm is implemented in a real-time diagnosis system for production microservice platforms.
  • RCD [13]: RCD is a scalable algorithm for detecting root causes of failures in complex microservice architectures using a hierarchical and localized learning approach. It treats the failure as an intervention to quickly identify the root cause, focuses learning on the relevant portion of the causal graph to avoid costly conditional independence tests, and explores the graph hierarchically. The technique is highly scalable, providing actionable insights about root causes, while traditional methods become infeasible due to high computation time.
  • BARO [39]: BARO is a robust, end-to-end RCA framework designed for multivariate time-series data in microservice systems. It integrates anomaly detection and root cause localization using Multivariate Bayesian Online Change Point Detection (BOCPD) to detect failures and estimate their occurrence times. For RCA, BARO introduces RobustScorer, a nonparametric hypothesis testing approach that ranks candidate root causes based on their distributional shifts, using median and interquartile range rather than mean and standard deviation to improve robustness.
  • Nezha [10]: Nezha is an interpretable and fine-grained root cause analysis (RCA) method for microservices that unifies heterogeneous observability data (metrics, traces, logs) into a homogeneous event format. This representation enables the construction of event graphs for integrated analysis. Nezha statistically localizes actionable root causes at granular levels, such as specific code regions or resource types, offering high interpretability to support confident mitigation actions by SREs.
Appendix F Figures for Clarity
We provide figures related to the system architecture and fault scenarios in this section, for better readability. The architecture of Product Review Platform is shown in Figure 5, and the system fault scenarios are demonstrated in Figure 6 and Figure 7.
Figure 5: Corresponding to Figure 1 (a). The architecture of Product Review Platform
Figure 6: Corresponding to Figure 3 left. Visualization of Cryptojacking system fault scenario. Right: External storage failure.
Figure 7: Corresponding to Figure 3 right. Visualization of External storage failure. system fault scenario.
Appendix G Dataset Labeling Methodology
We provide more details on the system fault labeling strategy, which comes in two-fold: the root cause labeling process and label validation.
Root Cause Labeling Process.
  • For each system fault, we designed controlled fault scenarios to mimic realistic fault patterns (e.g., external storage failure, database overload).
  • During each controlled fault case, we monitored system behaviors, including metrics and logs, to identify the exact root cause of the fault.
  • The ground truth root cause was then labeled based on the specific fault of the system. This ensures high accuracy in root cause labeling, as the faults were systematically induced and their impacts directly observed.
Label Validation.
  • To ensure label correctness, we validated the root cause labels by analyzing the system’s behavior during and after fault. This involved cross-checking the observed anomalies in system metrics and logs with the expected outcomes of the fault.
  • Multiple experts reviewed the labeled faults to confirm the consistency and correctness of the root cause assignments.
Appendix H Parameter Settings for All Baseline Models
We provide the detailed parameter settings for all baseline models as follows:
  • PC: alpha=0.05 (significance level for conditional independence tests), ci_test=’fisherz’ (type of conditional independence test).
  • RCD: ci_test=’chisq’ (type of conditional independence test). k=10 (top-k root causes), alpha_limit=0.5 (the maximum alpha for search),
  • $\varepsilon$-Diagnosis: test_size=0.6 (train test split ratio), root_cause_top_k=10 (top-k root causes),
  • Circa: test_size=0.6 (train test split ratio), root_cause_top_k=10 (top-k root causes),
  • Baro: Parameters of this algorithm are related to the Bayesian online change point detection. r=50 (magnitude of the hazard function for Bayesian online learning), k=3 (number of standard deviations from the mean to consider as an anomaly),
  • Nezha: level=service (detection at the service level)
Appendix I Dataset Representativeness
In this section, we aim to show the representativeness of the released dataset. While it is challenging to establish a universal metric for representativeness in benchmarks, we have made significant efforts to ensure the dataset covers diverse fault scenarios:
  • Real-World Fault Scenarios: The IT domain datasets (Product Review and Cloud Computing) encompass realistic microservice faults such as out-of-memory errors, DDoS attacks, and cryptojacking, as outlined in Section 3.1 and Appendix B. Similarly, the OT domain datasets (SWaT and WADI) include real-world cyber-physical system faults recorded in controlled environments.
  • Diversity of Fault Types: Across IT and OT domains, we include 10 distinct fault types, ensuring coverage of both transient and persistent system failures. This diversity reflects common issues faced by modern IT and OT systems.
  • Comparative Analysis: As seen in Table 3 and related discussions, our dataset exhibits performance trends consistent with other benchmarks (e.g., Petshop), supporting its credibility as a representative evaluation platform.
  • Quality Assurance: All data were collected using industry-standard monitoring tools like Prometheus, CloudWatch, and Elasticsearch. Each fault scenario was validated to ensure it mirrors real-world conditions.